# Generating Verifiable CoT from Execution-Traces

### v1

@@ Abstract.
We address this by grounding CoT generation directly in program execution traces. Our pipeline instruments code to capture its dynamic behavior, then narrates these verified execution traces into natural language rationales that are correct by construction. This execution-grounded approach ensures every reasoning step reflects what the program genuinely computes, eliminating logical hallucinations at the source. We evaluate our method on code reasoning tasks (forward reasoning on CruxEval and LiveCodeBench-Exec, backward reasoning on CruxEval-Input), as well as code generation and explanation tasks from HumanEval. Models trained on our bi-directional trace-grounded data achieve substantial improvements, with gains of up to 30 points on output prediction and 28 points on input prediction over base models, alongside improved explanation and code generation, demonstrating that verifiable reasoning fundamentally enhances model capabilities. https://github.com/shailja-thakur/Verified-Code-CoT
@@ 1. Introduction
Recent advances in large language models have enabled conversational code assistants that can generate, explain, and debug programs through natural language interaction. These systems—exemplified by GitHub Copilot, ChatGPT for code, and similar tools—have demonstrated remarkable syntactic fluency. However, they struggle with tasks requiring deep reasoning about program behavior: debugging requires understanding how state evolves, program comprehension demands tracing execution paths, and code refactoring needs verifying semantic preservation. These reasoning tasks are fundamental to software engineering practice, yet current models often produce plausible but incorrect explanations of program behavior.
@@
To bridge this gap, the research community has embraced fine-tuning models on Chain-of-Thought (CoT) data, which provides explicit, step-by-step rationales to guide the model’s reasoning process. However, the methods used to generate this data for code are often unsound. A powerful ”teacher” model generates the rationale, but these reasoning steps are not tethered to a program’s actual execution trace. The result is a dataset of ”logical hallucinations”—plausible but factually incorrect accounts of the program’s logic.
This challenge is especially acute in code reasoning, where small logical errors cascade into flawed understanding. Consider a model trained on a rationale claiming ”variable count is incremented in the loop”—when in fact, count is decremented. Such errors propagate during training, teaching the model to follow plausible but incorrect logical patterns. This lack of guaranteed soundness is a critical barrier for high-stakes applications. For debugging, a rationale that misrepresents variable state is actively misleading. For iterative self-refinement, feeding a model its own flawed reasoning causes compounding errors. For agentic workflows, a plan based on unsound understanding of a previous step can derail an entire task. Figure 1 illustrates this critical problem concretely.
Figure 1. Comparison of hallucinated vs trace-grounded CoT. Our approach translates pysnooper execution trace (showing line numbers, variable states, state transitions) into natural language reasoning. Color mapping shows trace sections: yellow=init, blue=iterations, green=return. LLM hallucinates wrong comparison (arr[2] vs arr[1] instead of arr[3]) and wrong search direction.
@@
Furthermore, effective code reasoning requires both forward inference (predicting outputs from inputs) and backward reasoning (understanding how outputs arose from inputs) (li2025codeio). This bi-directional capability is essential for debugging, code comprehension, and robust program analysis. However, current synthetic data generation methods do not systematically address this requirement, leading to models with asymmetric reasoning capabilities.
In this work, we bridge these gaps by introducing a new paradigm for data synthesis that achieves rationale-step verification at scale. We have developed a complete synthesis pipeline that generates CoT rationales by directly translating program execution traces into natural language. Specifically, we instrument a program to capture its dynamic execution state—recording variable values after each statement, control flow decisions, and state transitions derived from program semantic operations. This trace serves as an immutable ground truth from which rationales are derived. Each step in the resulting reasoning chain can be verified against the trace to confirm: (1) variable values match recorded runtime values, (2) state transitions correctly reflect how variables changed between statements, and (3) control flow accurately describes which branches executed. This ensures rationales are correct by construction, eliminating the logical hallucinations that plague other methods.
@@
    Chain of Thought Step Verification at Scale: We introduce the first execution-grounded synthesis pipeline that achieves rationale-step verification at scale.
@@
    Concept-First Curriculum Synthesis: A novel curriculum-driven generation approach that provides fine-grained control over problem complexity and algorithmic diversity, enabling systematic coverage from basic constructs to advanced algorithms.
@@
    Bi-Directional Reasoning Dataset: The first large-scale dataset (25,000 samples) systematically teaching both forward (input→output) and backward (output→input) reasoning with trace-grounded verification.
  4. (4)
    Comprehensive Empirical Validation: Through systematic ablations, we demonstrate that verification quality directly determines model performance.
@@ 2. Related Work
Verification Approaches in Code Generation. The term ”verification” appears frequently in code generation literature but often refers to different concepts. We identify three distinct approaches: (1) Outcome Verification validates only the final program output against expected results. This is the most common approach, used by benchmarks like HumanEval (chen2021evaluating) and methods like SemCoder (ding2024semcodertrainingcodelanguage) and rStar-Coder (liu2025rstarcoderscalingcompetitivecode). While effective for validating solutions, it offers no guarantee about the logical integrity of intermediate reasoning steps. (2) Formal Verification uses mathematical proofs (e.g., symbolic execution, SMT solvers) to guarantee program correctness. While sound, this approach is computationally heavyweight and impractical for large-scale LLM training data generation. (3) Rationale-Step Verification (our contribution) validates each reasoning step in the CoT chain. By generating rationales directly from a program’s execution trace—a factual record capturing variable values, state transitions, and control flow at runtime—we can verify that each reasoning step correctly describes: (a) the variable values at that point in execution, (b) how those values changed from the previous state due to program semantic operations, and (c) which control flow paths were taken. This verifies the logical integrity of the entire reasoning process, not just the final outcome.
@@
Recent work like TRACED (ding2024traced) and Execution Tuning (E.T.) (armengol2025execution) has explored using execution traces for pre-training, but these methods do not generate natural language rationales with step-by-step verification. Our work is the first to systematically achieve rationale-step verification at scale by translating execution traces into verifiable natural language CoT.
Comparison to SemCoder. A highly relevant work, SemCoder (ding2024semcodertrainingcodelanguage), also generates synthetic CoT data for fine-tuning. However, our methodology differs fundamentally in two aspects. First, SemCoder generates rationales via an LLM’s explanation of code behavior, then verifies only the final predicted output. We perform a direct translation of execution traces into natural language, making each intermediate reasoning step verifiable: claims about variable values can be checked against recorded runtime values, state transitions can be confirmed against logged changes between statements, and control flow descriptions can be validated against the actual execution path. This narration-vs-generation distinction eliminates logical hallucinations. Second, SemCoder uses a code-first approach seeded by existing programs. Our concept-first approach synthesizes problems from abstract concepts extracted from technical literature, providing fine-grained control over complexity and diversity. Conceptually closest to our trace translation, Jung et al.(jung2025codeexecutiongroundedsupervision) generate natural language CoT from execution traces, but focus on single-direction reasoning rather than our systematic bi-directional pipeline.
Comparison to rStar-Coder. rStar-Coder (liu2025rstarcoderscalingcompetitivecode) scales up verified data for competitive programming through iterative generation and execution-based verification. While both leverage execution, rStar-Coder verifies the final outcome of generated code, whereas we verify every intermediate reasoning step by generating rationales directly from traces. Consequently, rStar-Coder’s rationales are LLM-generated explanations (subject to hallucination), while ours are correct by construction. In summary, rStar-Coder excels at verifiably correct solutions; we create verifiably correct reasoning processes.
@@
As summarized in Table 1, no existing work provides a systematic pipeline for generating bi-directional CoT datasets where every reasoning step is grounded in verifiable execution. Our work addresses this gap, providing a trustworthy foundation for training robust code reasoning capabilities.
@@ 3. Data Synthesis Pipeline
Our data synthesis pipeline (Figure 2) is a multi-stage process designed to generate high-fidelity, bi-directional CoT data where every reasoning step is anchored in ground-truth program execution. This ensures that the resulting dataset is correct by construction, eliminating the logical hallucinations common in purely model-generated rationales.
@@ 3.1.1. Document Processing and Concept Extraction
Our pipeline processes a diverse corpus of permissively-licensed technical literature, including books from the StarCoder2-documentation dataset (starcoder2-doc) and curated programming resources spanning basic to advanced topics. Rather than relying on raw PDF text extraction, which produces noisy output, we employ Docling, a document understanding framework that renders PDFs into clean, structured markdown. This preprocessing preserves semantic structure (headings, code blocks, lists), removes pagination artifacts, and maintains proper formatting. The cleaned text is chunked into 4000-character segments with sliding overlap to prevent concept fragmentation.
@@ 3.1.4. Hierarchical Problem Synthesis
Step 1: Instruction Generation. Given a concept and its description, we prompt the teacher model to generate 6 distinct natural language problem instructions. The prompt is engineered for diversity along three axes: (1) problem domain (mathematics, finance, data processing, algorithms), (2) computational approach (iterative, recursive, dynamic programming, functional), and (3) difficulty (3 ”medium” targeting 30-60 LOC, 3 ”hard” targeting 50-100+ LOC). The prompt includes explicit anti-overlap guidance to avoid paraphrased versions of the same problem.
@@
Step 3: Code Generation. With the signature locked, we generate 5 candidate solutions per instruction by prompting with both the instruction and signature. The prompt explicitly instructs: (1) strictly adhere to the signature, (2) vary computational approach across implementations, and (3) ensure completeness.
@@
Step 5: Test Generation. We generate unit tests (3 test suites per instruction, each with up to 10 test functions) using a meticulously engineered prompt enforcing strict structural constraints. Each test is a top-level function (not a class method) containing exactly one assert statement with direct function calls using inline arguments—no variable assignments. Figure 5 illustrates the correct format versus prohibited patterns. These constraints are critical for Stage B’s granular pass/fail clustering and Stage C’s clean trace extraction, while enabling parallel sandboxed execution.
At the end of Stage A, we have for each concept: 6 instructions, each with 1 signature, 5 candidate solutions, and 3 test suites, totaling approximately 85,000 candidate instruction-solution-test triples.
@@ 3.2. Stage B: Execution-Based Verification and Agreement Clustering
The hierarchical generation in Stage A, while carefully constrained, inevitably produces noise: some candidate solutions are incorrect (due to LLM generation errors), some tests are malformed or have wrong expected outputs, and some instruction-signature pairs are ambiguous. Before generating expensive CoT rationales in Stage C, we must identify and discard low-quality artifacts. We adapt the Dual Agreement verification methodology from CodeT (chen2022codetcodegenerationgenerated), originally designed for single-solution selection, into a scalable batch filtering algorithm.
@@ 3.2.2. Dual Agreement Clustering
Best Cluster Selection. For each task, we rank clusters by score and select the single highest-scoring cluster. We extract one canonical solution from this cluster (selecting the shortest or most readable) and its associated passing tests. Algorithm 1 formalizes this procedure. This method effectively filters multiple noise sources: incorrect solutions land in low-scoring clusters due to failing many tests (low $|T_{p}|$), malformed tests reduce scores but cannot eliminate correct clusters if other valid tests exist, and ambiguous instructions cause solutions to fragment into multiple small clusters (low $|C_{i}|$), with the highest-scoring cluster representing the most common interpretation.
@@ 3.2.3. Theoretical Justification: Why Multiple Candidates Matter
As $K$ increases with sufficient diversity, $m$ (correct solutions) grows while incorrect solutions fragment into multiple small clusters with different error patterns, each with low $|C_{error}|$. This multiplicative scoring creates exponential separation, allowing correct clusters to dominate even with noisy candidates. However, this requires diversity: generating varied candidates via temperature sampling, prompt variation, and varied computational approaches in Stage A ensures that correct implementations converge while incorrect ones fragment. The effectiveness of this approach is empirically validated in Section LABEL:sec:dual_agreement_validation.
@@ 3.3.1. Step 1: Execution Trace Generation
For each verified solution and each of its passing test cases, we generate a ground-truth execution trace capturing the program’s dynamic behavior. We use pysnooper, a lightweight Python tracing library, to instrument the target function (identified via signature metadata from Stage A).
Instrumentation Strategy. Rather than tracing the entire program (which includes test framework overhead), we selectively instrument only the function under test. For a function solution, we decorate it with @pysnooper.snoop(output=trace_file), then execute the test case in a sandboxed subprocess with a strict timeout (10 seconds). This produces a line-by-line trace file detailing: (1) Call events: Function entry with input arguments, (2) Line execution: Each line number executed in sequence, (3) Variable modifications: Every assignment with variable name, old value, and new value (capturing state transitions), and (4) Return events: Function exit with return value.
@@ 3.3.5. Step 5: Final Assembly into Conversational Format
This yields our final dataset of 54k verified, trace-grounded, bi-directional reasoning samples.
@@ 4.2.1. Hyperparameter Selection for Data Generation
    Difficulty-Based Subsetting: From the master dataset, we created three distinct subsets for our ablation studies, based on different filtering strategies:
@@
    • Model-Derived Difficulty Set (25k): This high-quality subset was created via a novel filtering strategy: we kept only the problems from the 54k set that an LLM failed to solve correctly. This creates a targeted curriculum focused on the model’s specific weaknesses.
@@ 4.3. Ablation Study Plan
Our experiments are designed as a sequential, top-down filter to efficiently identify the optimal data configuration. While testing every combination of our data subsets and reasoning formats would be ideal, such a full factorial experiment is computationally infeasible. Our sequential approach is a practical and methodologically sound alternative that allows us to isolate the impact of our two primary contributions: difficulty-filtered subsets and CoT directionality. We operate on the reasonable assumption that the data subset demonstrating the highest quality for forward CoT will also be the most effective foundation for the other CoT formats. All experiments are conducted in a Supervised Fine-Tuning (SFT) setting.

### v2

# Generating Verifiable Chain of Thoughts from Exection-Traces

@@ Abstract.
We address this by grounding CoT generation directly in program execution traces. Our pipeline instruments code to capture its dynamic behavior, then narrates these execution traces into natural language and factually-grounded rationales that are verifiable by design. This execution-grounded approach ensures every reasoning step reflects what the program computes, eliminating logical hallucinations at the source. We evaluate our method on code reasoning tasks, code generation and explanation tasks from HumanEval. Models trained on our bi-directional trace-grounded data achieve substantial improvements on reasoning tasks, with gains of up to 30 points on output prediction and 28 points on input prediction over base models, alongside competitive explanation and code generation performance. https://github.com/shailja-thakur/Verified-Code-CoT
@@ 1. Introduction
Recent advances in large language models have enabled code assistants that can generate, explain, and debug programs through natural language interaction. These systems—exemplified by GitHub Copilot, Claude Code, and similar tools—have demonstrated remarkable syntactic fluency. However, they struggle with tasks requiring deep reasoning about program behavior: debugging requires understanding how state evolves, program comprehension demands tracing execution paths, and code refactoring needs verifying semantic preservation. These reasoning tasks are fundamental to software engineering practice, yet current models often produce plausible but incorrect explanations of program behavior.
@@
To bridge this gap, the research community has embraced fine-tuning models on Chain-of-Thought (CoT) data, which provides explicit, step-by-step rationales to guide the model’s reasoning process. However, the methods used to generate this data for code are often unsound. A powerful teacher model generates the rationale, but these reasoning steps are not tethered to a program’s actual execution trace. The result is a dataset of logical hallucinations—plausible but factually incorrect accounts of the program’s logic.
This challenge is especially acute in code reasoning, where small logical errors cascade into flawed understanding. Consider a model trained on a rationale claiming ”variable count is incremented in the loop”—when in fact, count is decremented. Such errors propagate during training, teaching the model to follow plausible but incorrect logical patterns. This lack of guaranteed soundness is a critical barrier for high-stakes applications. For debugging, a rationale that misrepresents variable state is misleading. For iterative self-refinement, feeding a model its own flawed reasoning causes compounding errors. For agentic workflows, a plan based on unsound understanding of a previous step can derail an entire task. Figure 1 illustrates this critical problem concretely.
Figure 1. Comparison of hallucinated vs trace-grounded CoT. Our approach translates pysnooper (pysnooper) execution trace (showing line numbers, variable states, state transitions) into natural language reasoning. Color mapping shows trace sections: yellow=init, blue=iterations, green=return. LLM hallucinates wrong comparison (arr[2] vs arr[1] instead of arr[3]) and wrong search direction.
@@
Furthermore, effective code reasoning requires both forward inference (predicting outputs from inputs) and backward reasoning (understanding how outputs arose from inputs) (li2025codeio). This bi-directional capability is essential for debugging, code comprehension, and program analysis. However, current synthetic data generation methods do not systematically address this requirement, leading to models with asymmetric reasoning capabilities.
In this work, we bridge these gaps by introducing a new paradigm for data synthesis that achieves rationale-step verification at scale. We have developed a complete synthesis pipeline that generates CoT rationales by directly translating program execution traces into natural language. Specifically, we instrument a program to capture its dynamic execution state—recording variable values after each statement, control flow decisions, and state transitions derived from program semantic operations. This trace serves as an ground truth from which rationales are derived. Each step in the resulting reasoning chain can be verified against the trace to confirm: (1) variable values match recorded runtime values, (2) state transitions correctly reflect how variables changed between statements, and (3) control flow accurately describes which branches executed. This ensures rationales are verifiable by design, eliminating the logical hallucinations that plague other methods.
@@
    Chain of Thought Step Verification at Scale: We introduce a scalable execution-grounded synthesis pipeline that achieves rationale-step verification.
@@
    Concept-First Curriculum Synthesis: A curriculum-driven synthesis that provides control over problem complexity and diversity, enabling systematic coverage from basic constructs to advanced algorithms.
@@
    Comprehensive Empirical Validation: Through systematic validations, we demonstrate that execution-groundedness and bi-directional nature of CoTs directly determine model performance.
@@ 2. Related Work
Verification Approaches in Code Generation. The term ”verification” in code generation literature often refers to different concepts. We identify three distinct approaches:
  1. (1)
    Outcome Verification validates only the final program output against expected results. This is the most common approach, used by benchmarks like HumanEval (chen2021evaluating) and methods like SemCoder (ding2024semcodertrainingcodelanguage) and rStar-Coder (liu2025rstarcoderscalingcompetitivecode). While effective for validating solutions, it offers no guarantee about the logical integrity of intermediate reasoning steps.
  2. (2)
    Formal Verification uses mathematical proofs (e.g., symbolic execution, SMT solvers) to guarantee program correctness. While sound, this approach is computationally heavyweight and impractical for large-scale training data generation.
  3. (3)
    Rationale-Step Verification (our contribution) validates each reasoning step in the CoT chain. By generating rationales directly from a program’s execution trace—a factual record capturing variable values, state transitions, and control flow—we verify that each reasoning step correctly describes runtime states and semantic operations. This verifies the logical integrity of the entire reasoning process, not just the outcome.
@@
Recent work like TRACED (ding2024traced) and Execution Tuning (E.T.) (armengol2025execution) has explored using execution traces for pre-training, but these methods do not generate natural language rationales with step-by-step verification. Our work is the first to systematically achieve rationale-step verification at scale by translating execution traces into verifiable natural language CoT. Conceptually closest to our trace translation, Jung et al. generate natural language CoT from execution traces but focus on single-direction reasoning rather than a systematic bi-directional pipeline.
Comparison to SemCoder and rStar-Coder. SemCoder and rStar-Coder also generate synthetic CoT data, but our methodology differs in two aspects. First, while SemCoder generates rationales via an LLM’s explanation and verifies only output, we perform a direct translation of execution traces, eliminating logical hallucinations. Second, unlike SemCoder’s code-first approach, our concept-first approach provides fine-grained control over complexity and diversity by synthesizing problems from abstract technical concepts. Similarly, rStar-Coder excels at verifiably correct solutions, while we focus on creating verifiably correct reasoning processes.
@@
As summarized in Table 1, no existing work provides a systematic pipeline for generating bi-directional CoT datasets where every reasoning step is grounded in verifiable execution. Our work addresses this gap, providing a foundation for training robust code reasoning capabilities.
@@ 3. Data Synthesis Pipeline
Our data synthesis pipeline (Figure 2) is a multi-stage process designed to generate high-fidelity, bi-directional CoT data where every reasoning step is anchored in ground-truth program execution.
@@ 3.1.1. Document Processing and Concept Extraction
Our pipeline processes a diverse corpus of permissively-licensed technical literature, including books from the StarCoder2-documentation dataset (starcoder2-doc) and curated programming resources spanning basic to advanced topics. Rather than relying on raw PDF text extraction, which produces noisy output, we employ Docling (Docling), a document understanding framework that renders PDFs into clean, structured markdown. This preprocessing preserves semantic structure (headings, code blocks, lists), removes pagination artifacts, and maintains proper formatting. The cleaned text is chunked into 4000-character segments with sliding overlap to prevent concept fragmentation.
@@ 3.1.4. Hierarchical Problem Synthesis
Step 1: Instruction Generation. Given a concept and its description, we prompt the teacher model to generate distinct natural language problem instructions. The prompt is engineered for diversity along three axes: (1) problem domain (mathematics, finance, data processing, algorithms), (2) computational approach (iterative, recursive, dynamic programming, functional), and (3) difficulty (”medium” targeting 30-60 LOC, ”hard” targeting 50-100+ LOC). The prompt includes explicit anti-overlap guidance to avoid paraphrased versions of the same problem.
@@
Step 3: Code Generation. With the signature locked, we generate candidate solutions per instruction by prompting with both the instruction and signature. The prompt explicitly instructs: (1) strictly adhere to the signature, (2) vary computational approach across implementations, and (3) ensure completeness.
@@
Step 5: Test Generation. We generate unit tests (3 test suites per instruction, each with up to 10 test functions) using a meticulously engineered prompt enforcing strict structural constraints. Each test is a top-level function (not a class method) containing exactly one assert statement with direct function calls using inline arguments—no variable assignments. Figure 5 illustrates the correct format versus prohibited patterns. These constraints are critical for Stage B granular pass/fail clustering and Stage C clean trace extraction, while enabling parallel sandboxed execution.
@@ 3.2. Stage B: Execution-Based Verification and Agreement Clustering
The hierarchical generation in Stage A, while carefully constrained, produces noise: some candidate solutions are incorrect (due to LLM generation errors), some tests are malformed or have wrong expected outputs, and some instruction-signature pairs are ambiguous. Before generating expensive CoT rationales in Stage C, we must identify and discard low-quality artifacts. We adapt the Dual Agreement verification methodology from CodeT (chen2022codetcodegenerationgenerated), originally designed for single-solution selection, into a scalable batch filtering algorithm.
@@ 3.2.2. Dual Agreement Clustering
Best Cluster Selection. For each task, we rank clusters by score and select the single highest-scoring cluster. We extract one canonical solution from this cluster (selecting the pair with maximum code coverage) and its associated passing tests. Algorithm 1 formalizes this procedure. This method effectively filters multiple noise sources: incorrect solutions land in low-scoring clusters due to failing many tests (low $|T_{p}|$), malformed tests reduce scores but cannot eliminate correct clusters if other valid tests exist, and ambiguous instructions cause solutions to fragment into multiple small clusters (low $|C_{i}|$), with the highest-scoring cluster representing the most common interpretation.
@@ 3.2.3. Theoretical Justification: Why Multiple Candidates Matter
As $K$ increases with sufficient diversity, $m$ (correct solutions) grows while incorrect solutions fragment into multiple small clusters with different error patterns, each with low $|C_{error}|$. This multiplicative scoring creates exponential separation, allowing correct clusters to dominate even with noisy candidates. However, this requires diversity: generating varied candidates via temperature sampling, prompt variation, and varied computational approaches in Stage A ensures that correct implementations converge while incorrect ones fragment. The effectiveness of this approach is empirically validated in Section 5.1.
@@ 3.3.1. Step 1: Execution Trace Generation
For each verified solution and selected passing test cases, we generate a ground-truth execution trace capturing the program’s dynamic behavior. We use pysnooper(pysnooper), a lightweight Python tracing library, to instrument the target function (identified via signature metadata from Stage A).
Instrumentation Strategy. Rather than tracing the entire program (which includes test framework overhead), we selectively instrument only the function under test. This approach is robust against issues like recursion or complex interprocedural calls because the structural constraints imposed during Code Generation (Stage A)—which enforce a flattened logic without nested functions—guarantee that all core computation resides within the instrumented scope. For a function solution, we decorate it with @pysnooper.snoop(output=trace_file), then execute the test case in a sandboxed subprocess with a strict timeout (10 seconds). This produces a line-by-line trace file detailing: (1) Call events: Function entry with input arguments, (2) Line execution: Each line number executed in sequence, (3) Variable modifications: Every assignment with variable name, old value, and new value (capturing state transitions), and (4) Return events: Function exit with return value.
@@ 4.2.1. Hyperparameter Selection for Data Generation
    Difficulty-Based Subsetting: From the master dataset of 85k code-test pairs, we created three distinct subsets for our ablation studies, based on different filtering strategies:
@@
    • Model Answerability Set (25k): This high-quality subset was created via a novel filtering strategy: we kept only the problems from the 54k deduplicated set that an LLM failed to solve correctly. This creates a targeted curriculum focused on the model’s specific weaknesses.
@@ 4.3. Ablation Study Plan
Our experiments are designed as a sequential, top-down filter to efficiently identify the data configuration. While testing every combination of our data subsets and reasoning formats would be ideal, such a full factorial experiment is computationally infeasible. Our sequential approach is a practical and methodologically sound alternative that allows us to isolate the impact of our two primary contributions: difficulty-filtered subsets and CoT directionality. We operate on the reasonable assumption that the data subset demonstrating the highest quality for forward CoT will also be the most effective foundation for the other CoT formats. All experiments are conducted in a Supervised Fine-Tuning (SFT) setting.